Circular No. 8286

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 C1 (LARSEN) J. A. Larsen, Lunar and Planetary Laboratory, reports his discovery of

a comet on Spacewatch CCD images obtained with the 0.9-m f/3 reflector on Feb. 12.42 UT (discovery observation below). Following posting on the NEO Confirmation Page, cometary appearance was also reported on CCD images taken around Feb. 13.3 by J. Young (Table Mountain, 0.6-m reflector; 6′′ coma of total mag 18.0 with little or no central condensation and a short, stubby tail 14′′ long in p.a. 260^{o}) and by J. E. McGaha (Tucson, AZ, 0.36-m reflector; eighteen stacked images show a 6′′ coma and a 5′′ fanshaped tail in p.a. 90^{o}).

2004 UT  α_{2000}  δ_{2000}  Mag.
Feb. 12.42045  10^{h}16^{m}47.^{s}43  +9^{o}41′55′′.1  20.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-C52.

T = 2005 Oct. 12.503 TT  ω = 70.^{o}356
Ω = 151.630 }2000.0
q = 4.26365 AU  i = 56.251

SUPERNOVAE 2004T, 2004U, AND 2004W Further to IAUC 8277, M. Moore and W. Li report the LOSS discovery

of a supernova, designated SN 2004W and located at α = 12^{h}43^{m}36^{s}.52, δ = +11^{o}31′50′′.8 (equinox 2000.0), which is 51′′.6 west and 78′′.7 south of the nucleus of NGC 4649. Approximate KAIT magnitudes: 2003 June 4.2 UT, [20.0; 2004 Jan. 28.5, 18.8; 29.5, 18.9; Feb. 11.5, 19.3. A. V. Filippenko, R. Chornock, and R. J. Foley, University of California at Berkeley, report that inspection of CCD spectra (range 310–920 nm), obtained on Feb. 13 with the Keck I 10-m telescope (+ LRIS), shows that SN 2004W is of type Ia, specifically the underluminous SN 1991bg variety (Filippenko et al. 1992, A.J. 104, 1543; Ruiz-Lapuente et al. 1993, Nature 365, 728), roughly half a year past maximum brightness. The spectrum is dominated by very strong, broad [Ca II] 730-nm emission, with weaker iron emission lines at blue and visual wavelengths, as well as the Ca II near-infrared triplet.

Corrigenda. On IAUC 8285, the offset for SN 2004T should read 6′′.1 E, 10′′.7 N (not south). The declination for SN 2004U should read +28^{o}32′23′′.6 (not 22′).

2004 February 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green